OpenStudy (anonymous):

verify. cosxtan^3x=sinxtan^2x

OpenStudy (anonymous):

u have to use\[\tan x=\frac{\sin x}{\cos x}\]

OpenStudy (anonymous):

yeahh but im kind of confused on how to for this questions

OpenStudy (anonymous):

\(\cos x\tan^3x=\cos x \tan x \ \tan^2x\) ------

OpenStudy (anonymous):

how did you get cosxtanx for the right side?

OpenStudy (anonymous):

well exponent of \(\tan x\) is 3 i can write it like this\[\tan^3 x=\tan x \tan^2 x\]

OpenStudy (anonymous):

like \(a^3=a.a^2\)

OpenStudy (anonymous):

oh okayy

OpenStudy (anonymous):

um and then what is the next step ?

OpenStudy (ajprincess):